The following budget data pertain to the Machining Department of Yolkenverst Co.:

Maximum capacity 60,000 units
Machine hours per unit 2.50
Variable factory overhead $3.60 per machine hour
Fixed factory overhead $433,500

The company prepared the budget at 85% of the maximum capacity level. The department uses machine hours as the basis for applying standard factory overhead costs to production.

During the year the Machining Department produced 50,000 units, consuming 127,500 machine hours and incurring $433,500 of fixed overhead. For the current year the department has a fixed overhead production volume variance, rounded to the nearest whole dollar, of:
Answer #1

Solution:

Budgted Hours = 60000*85%* 2.5 = 127,500

Fixed Overhead application rate = $433500 / 127500 = $3.40 per machine hour

Fixed overhead applied = 50000*2.50*$3.40 = $425,000

fixed overhead production volume variance = Applied Fixed Overehad - Budgeted Fixed Overhead

= $425000 - $433500

= $8500 Unfavorable